OpenStudy (anonymous):

Three people A,B and C, share 416 dollar among themselves.A receives 169 dollar and B receives 156 dollar.find the ratio in which the sum of money is shared.

OpenStudy (anonymous):

@Loser66 @AlexandervonHumboldt2 @conain007 @xapproachesinfinity @AravindG

OpenStudy (anonymous):

@iamabarbiegirl @undeadknight26

OpenStudy (loser66):

169+156=?

OpenStudy (anonymous):

325 ....=169+156

OpenStudy (loser66):

and \(\dfrac{325}{416}=??\)

OpenStudy (anonymous):

y we will divide 325/416???

OpenStudy (loser66):

sure

OpenStudy (anonymous):